# Resolving Turbulent Magnetohydrodynamics: A Hybrid Operator-Diffusion Framework

### v1

@@ 1 Introduction
Magnetized fluid turbulence—characterized by intricate coupling between fluid and magnetic fields—remains one of the most formidable challenges in theoretical and computational physics. The two-dimensional, incompressible, resistive magnetofluid, defined by the magnetohydrodynamic (MHD) equations, hosts a rich tapestry of phenomena: Lorentz-force–driven advection, resistive diffusion, current-sheet formation, magnetic reconnection, and energetic cascades spanning scales. Even in this simplified two-dimensional, incompressible regime, the dynamics far exceed the complexity of traditional testbeds such as Kolmogorov flow, 2D Navier–Stokes turbulence in vorticity form, steady-state Darcy flow, Burgers equations in the scalar, inviscid and vector types, and the 2D linear and nonlinear shallow water equations [1, 2].
@@
The respective limitations of PINOs and diffusion models suggest a natural synergy. This motivates the adoption of a hybrid framework, such as the one proposed by Oommen et al. [14], which synergistically fuses the strengths of both paradigms. In this two-stage approach, a deterministic PINO provides the foundational spatio-temporal structure and enforces physical consistency, while a stochastic, score-based diffusion model acts as a corrector, generating the missing high-frequency turbulent details. While this methodology has shown promise, its application to a physically rich and challenging system like magnetized turbulence remains unexplored. Through this work, we harness the PINO’s operator-level structure and the diffusion model’s fine-scale generation capabilities—arguably reaching the necessary nuance to capture full-field statistics, coherent structures, and non-Gaussian tails in turbulence up to $\mathrm{Re}\approx 3{,}000$.
@@
In this Letter, we present the first PINO–Diffusion hybrid tailored for turbulent MHD. We demonstrate that it achieves unprecedented fidelity: matching spectral distributions, recovering intermittent phenomena, and generalizing across forcing and magnetic Prandtl numbers—significantly outperforming deterministic surrogate models. Our method not only advances the frontier of AI-driven turbulence modeling but lays the foundation for broader applications in more complex, physically rich MHD systems.
2 Methodology
@@ 2.1 Governing Equations and Data Generation
subject to the constraints $\nabla\cdot\mathbf{u}=0$ and $\nabla\cdot\mathbf{B}=0$. Here, $\mathbf{u}$ is the fluid velocity, $\mathbf{B}$ is the magnetic field, $p$ is the pressure, $\rho_{0}=1$ is the constant fluid density, $\nu$ is the kinematic viscosity, and $\eta$ is the magnetic diffusivity. The kinetic and magnetic Reynolds numbers are defined as $\mathrm{Re}=1/\nu$ and $\mathrm{Re_{m}}=1/\eta$, respectively. For this work, we consider a magnetic Prandtl number $\mathrm{Pr_{m}}=\nu/\eta=1$, such that $\mathrm{Re}=\mathrm{Re_{m}}$.
@@
This system of equations was solved numerically using the open-source spectral solver Dedalus [16], closely following the numerical framework established in Ref. [7]. The simulations were configured on a doubly periodic 2D domain $[0,1]^{2}$, which was discretized using a Fourier basis with a primary resolution of $128\times 128$ grid points and a de-aliasing factor of $3/2$. A fourth-order Runge-Kutta scheme was employed for time integration with a fixed timestep of $\Delta t=10^{-3}$, evolving the system to a final time of $t=1.0$ with snapshots saved every $0.01$ time units. This process yields 101 high-resolution temporal frames for each simulation, which were then temporally subsampled by a factor of 4 to produce the final 26-frame sequences used for training and evaluation. To create a diverse ensemble of scenarios, initial conditions were generated from Gaussian Random Fields (GRFs) [17] with a characteristic length scale of $l=0.1$, ensuring the initial velocity and magnetic fields were divergence-free. To investigate the transition from laminar to turbulent flow, we generated distinct datasets for a wide range of Reynolds numbers, specifically $\mathrm{Re}\in\{100,250,500,750,1000,3000,10000\}$. The generation of this extensive dataset was performed on the NCSA Delta cluster, where we employed a massively parallel approach to distribute thousands of independent simulation instances using all 128 CPU cores of a compute node. For each Reynolds number, this process yielded an ensemble of 1000 simulations, which was randomly partitioned into a training set of 800 samples, a validation set of 100 samples, and a test set of 100 samples.
@@ 2.2 Diffusion-Enhanced Operator Framework
The first stage of our pipeline employs a Physics-Informed Tensor-Factorized Fourier Neural Operator (PINO) [4, 1] to learn the mean-field evolution of the MHD system. The architecture, an extension of the model used in Ref. [7], is designed to process 3D spatio-temporal data ($x,y,t$). It consists of 8 Fourier layers with 32 latent channels. Within each spectral convolution, we retain 8 Fourier modes and utilize a Canonical Polyadic (CP) tensor decomposition with a rank of 0.5 to factorize the spectral weights [18, 19]. This factorization enhances model expressivity while maintaining computational tractability on large-scale hardware. Crucially, the PINO is trained with a composite loss function that, in addition to a standard data-fitting term, includes residuals of the governing MHD equations (Eqs. 1-3) as soft constraints. The PINO was trained on a single AMD MI250x GPU on the Frontier supercomputer at the Oak Ridge Leadership Computing Facility (OLCF).
The output of the trained PINO, which represents a smooth, spectrally-limited approximation of the true solution, serves as a conditional prior for the second stage: a score-based diffusion model [9]. This generative model is built upon a UNet backbone [20] with a base dimension of 128 and six down/up-sampling stages, defined by channel multipliers of (1, 2, 3, 5, 8, 12). Self-attention, accelerated via Flash Attention [21, 22], is incorporated at each resolution with 8 heads and a head dimension of 64. This UNet is integrated into the elucidated diffusion framework of Karras et al. [23], which employs a principled set of preconditioning functions and a specific noise schedule to normalize the network’s learning task, leading to more effective training and ultimately higher-quality sample generation. The model is tasked with stochastically generating the high-wavenumber turbulent structures absent from the PINO’s prediction. During its reverse-denoising process, it is explicitly conditioned on the PINO output, ensuring that the generated turbulent structures are physically consistent with the underlying large-scale flow. The training data for the diffusion model, consisting of pairs of PINO predictions and their corresponding ground truth simulation states, was generated on Frontier, while the model itself was trained on NVIDIA H100 GPUs on the NCSA Delta AI cluster.
@@ 3.2 Spectral Correction via Diffusion-Enhanced Framework
The generative power of this corrector stage stems from its training via denoising score matching [24], where it learns to approximate the score, or the gradient of the log-probability of the data, $\nabla_{x}\log p(x;\sigma)$, at varying noise levels $\sigma$. As detailed by Karras et al. [23], inference is then framed as solving a reverse-time stochastic differential equation (SDE). This process iteratively evolves a random noise field towards a clean sample by following the learned score function while simultaneously injecting a controlled amount of new noise at each step. This stochasticity helps to correct for accumulated errors and improves sample quality.
@@ 4 Summary and Conclusions
In this Letter, we addressed the challenge of constructing accurate surrogate models for two-dimensional, incompressible MHD, particularly in regimes characterized by complex, multi-scale turbulence. While PINOs are effective in capturing large-scale, equation-constrained dynamics, we demonstrated that they are fundamentally limited by an inherent spectral bias that suppresses high-wavenumber content [7]. This limitation becomes especially pronounced at moderate to high Reynolds numbers, where fine-scale magnetic and velocity structures dominate the system evolution.
@@
These findings establish a promising paradigm for surrogate modeling in computational physics. By delegating the learning of low-frequency, deterministic evolution to a neural operator and the high-frequency, stochastic corrections to a generative model, the proposed framework provides a principled way to overcome the spectral limitations of existing approaches. This divide-and-conquer strategy is likely to generalize to a broad class of multi-scale problems in fluid dynamics, astrophysics, and plasma physics.
Future work may explore infusing physics-based constraints—such as PDE residuals—into the diffusion stage to enhance physical consistency [10], as well as extending the framework to kinetic plasma models. These directions offer compelling opportunities to further improve fidelity, generalizability, and interpretability in the data-driven modeling of complex physical systems.
@@ Acknowledgements
EAH acknowledges the partial support of NSF grants OAC-2514142 and OAC-2209892. RH acknowledges partial support from NSF grants OAC-2004879, OAC-2005572, OAC-2103680, OAC-2310548, OAC-2411068. RH acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used supporting resources at the Argonne and the Oak Ridge Leadership Computing Facilities. The Argonne Leadership Computing Facility at Argonne National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357. The Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC05-00OR22725. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
@@
The complete source code for the Diffusion-Integrated Neural Operators (DINOs) framework developed for this study, including all modules for data generation, model training, and analysis, is publicly available in the official repository at https://github.com/semihkacmaz/DINOs.
Appendix A Technical Background
A.1 Neural Operators and Physics-Informed Training
Definition A.1.1 (Neural Operator).
Definition A.1.2 (Kernel Integral Operator).
[…]
Definition A.1.3 (Fourier Integral Operator).
The Fourier integral operator $\mathcal{K}_{\phi}$ is defined as:
The training of this operator is guided by a composite, physics-informed loss function. The total loss, $\mathcal{L}(\theta)$, for a set of model parameters $\theta$ is a weighted sum of several components:
The Data Loss ($\mathcal{L}_{data}$) is a standard supervised term measuring the discrepancy between the model’s prediction and the ground truth. The Initial Condition Loss ($\mathcal{L}_{ic}$) specifically enforces this agreement at $t=0$. The crucial PDE Residual Loss ($\mathcal{L}_{pde}$) embeds the governing physics by penalizing the norm of the residual when the model’s output is passed through the known differential operators (Eqs. 1-3). This composite objective, inspired by Physics-Informed Neural Networks (PINNs) [3], guides the optimization towards solutions that are both accurate and physically consistent.
A.2 Score-Based Diffusion Models and Conditional Generation
The generative model used in the second stage of our framework is a score-based diffusion model. This class of models is built on the principle of reversing a continuous-time stochastic differential equation (SDE) that transforms a data distribution, $p_{data}(x)$, into a simple prior distribution [9]. A forward SDE gradually perturbs a data sample $x_{0}$ into a noise vector $x_{T}$:
$dx=f(x,t)dt+g(t)dw.$ (9)
$dx=[f(x,t)-g(t)^{2}\nabla_{x}\log p_{t}(x_{t})]dt+g(t)d\bar{w},$ (10)
There are two primary strategies to implement this in practice. One approach is to train an unconditional score model $s_{\theta}(x_{t},t)\approx\nabla_{x}\log p_{t}(x_{t})$ and then, during inference, separately estimate the guidance term, often via the gradient of a separately trained classifier. This technique, known as classifier guidance, was a key innovation for improving conditional sample quality [26].
However, the method used in our work, which is common in modern diffusion model implementations, takes a more direct approach that builds upon the principles of Denoising Diffusion Probabilistic Models (DDPMs) [8]. We train a single neural network to directly approximate the full conditional score, $s_{\theta}(x_{t},y,t)\approx\nabla_{x}\log p_{t}(x_{t}|y)$. The conditioning information $y$ (the PINO output) is provided to the network as an additional input, typically by concatenating the condition $y$ with the noisy state $x_{t}$ along the channel dimension. The network is then trained to denoise this composite input.
The efficacy of this conditional framework hinges on a crucial division of labor. The score-based denoising process is inherently more effective at refining high-wavenumber components than low-wavenumber ones, where the score function gradient is weaker. Consequently, for the framework to succeed, the neural operator must provide an accurate low-frequency prior, or ‘mean flow’. A formal justification for this requirement, which analyzes the behavior of the score function in the frequency domain, is provided in the appendix of Oommen et al. [14].
Appendix B Detailed Results for the High-Turbulence Regime

### v2

@@ 1 Introduction
Magnetized fluid turbulence—characterized by intricate coupling between fluid and magnetic fields—remains one of the most formidable challenges in theoretical and computational physics. Beyond its intrinsic theoretical appeal, the study of magnetohydrodynamic (MHD) turbulence carries significant scientific and practical importance. In astrophysics, MHD turbulence underpins the transport of angular momentum in accretion disks [1, 2], governs the heating of the solar corona [3, 4], and regulates the structure of the interstellar medium [5, 6]. In geophysics, it plays a central role in the generation and sustenance of planetary dynamos, including Earth’s magnetic field [7]. In laboratory and engineering contexts, understanding MHD turbulence is crucial for optimizing magnetic confinement in fusion devices such as tokamaks and stellarators [8, 9], as well as for liquid-metal cooling technologies in next-generation nuclear reactors [10]. Across these systems, nonlinear couplings between velocity and magnetic fields govern cascade processes, intermittency, and reconnection events that directly impact energy transport, stability, and confinement. Thus, accurate modeling of MHD turbulence is indispensable not only for advancing fundamental plasma physics but also for enabling predictive capability in astrophysical, geophysical, and technological applications.
The two-dimensional, incompressible, resistive magnetofluid, defined by the magnetohydrodynamic (MHD) equations, hosts a rich tapestry of phenomena: Lorentz-force–driven advection, resistive diffusion, current-sheet formation, magnetic reconnection, and energetic cascades spanning scales. Even in this simplified two-dimensional, incompressible regime, the dynamics far exceed the complexity of traditional testbeds such as Kolmogorov flow, 2D Navier–Stokes turbulence in vorticity form, steady-state Darcy flow, Burgers equations in the scalar, inviscid and vector types, and the 2D linear and nonlinear shallow water equations [11, 12].
@@
The respective limitations of neural operators and diffusion models suggest a natural synergy. This motivates the adoption of a hybrid framework, such as the one proposed by Oommen et al. [24], which synergistically fuses the strengths of both paradigms. In this two-stage approach, a deterministic neural operator provides the foundational spatio-temporal structure and enforces physical consistency, while a stochastic, score-based diffusion model acts as a corrector, generating the missing high-frequency turbulent details. While this methodology has shown promise, its application to a physically rich and challenging system like magnetized turbulence remains unexplored. Through this work, we harness the PINO’s operator-level structure and the diffusion model’s fine-scale generation capabilities—arguably reaching the necessary nuance to capture full-field statistics, coherent structures, and non-Gaussian tails in turbulence up to $\mathrm{Re}\approx 3{,}000$.
@@
Here, we present the first PINO–Diffusion hybrid tailored for turbulent MHD. We demonstrate that it achieves unprecedented fidelity: matching spectral distributions, recovering intermittent phenomena, and significantly outperforming deterministic surrogate models. Our method not only advances the frontier of AI-driven turbulence modeling but lays the foundation for broader applications in more complex, physically rich MHD systems.
2 Background and Methodology
This section outlines the theoretical foundations of the neural operator and score-based diffusion models. It further describes how these components were implemented and integrated for the specific physical problem addressed in this work.
2.1 Neural Operators and Physics-Informed Training
Definition 2.1.1 (Neural Operator).
Definition 2.1.2 (Kernel Integral Operator).
$(\mathcal{K}(a;\phi)v_{t})(x)=\int_{D}\kappa_{\phi}(x,y,a(x),a(y))v_{t}(y)dy,$ (2)
Definition 2.1.3 (Fourier Integral Operator).
The Fourier integral operator $\mathcal{K}_{\phi}$ is defined as:
To ensure physical consistency and data fidelity, the operator is trained using a composite physics-informed loss function. The total loss, $\mathcal{L}(\theta)$, for a set of model parameters $\theta$ is a weighted sum of several components:
The Data Loss ($\mathcal{L}_{data}$) is a standard supervised term measuring the discrepancy between the model’s prediction and the ground truth. The Initial Condition Loss ($\mathcal{L}_{ic}$) specifically enforces this agreement at $t=0$. The crucial PDE Residual Loss ($\mathcal{L}_{pde}$) embeds the governing physics by penalizing the norm of the residual when the model’s output is passed through the known differential operators (Eqs. 9-11). This composite objective, inspired by Physics-Informed Neural Networks (PINNs) [13], guides the optimization towards solutions that are both accurate and physically consistent.
2.2 Score-Based Diffusion Models and Conditional Generation
The second stage of our framework acts as a corrector: a score-based diffusion model conditioned on the low-frequency output of the trained PINO. This generative component enhances the surrogate solution by recovering fine-scale structures omitted in the operator’s smooth approximation.
This class of models is built on the principle of reversing a continuous-time stochastic differential equation (SDE) that transforms a data distribution, $p_{data}(x)$, into a simple prior distribution [19]. A forward SDE gradually perturbs a data sample $x_{0}$ into a noise vector $x_{T}$:
$dx=f(x,t)dt+g(t)dw.$ (5)
$dx=[f(x,t)-g(t)^{2}\nabla_{x}\log p_{t}(x_{t})]dt+g(t)d\bar{w},$ (6)
There are two primary strategies to implement conditional generation with diffusion models. One approach is to train an unconditional score model $s_{\theta}(x_{t},t)\approx\nabla_{x}\log p_{t}(x_{t})$, and then, during inference, modify the reverse diffusion process by incorporating a guidance term derived from an external classifier. This technique, known as classifier guidance, involves using the gradient of the log-probability from a separately trained classifier $\nabla_{x}\log p(y|x_{t})$ to steer the sampling process toward samples that are more likely to belong to a desired class $y$. Classifier guidance was a key innovation for improving the fidelity and class-conditional accuracy of generated samples [26].
However, the method used in our work, which is common in modern diffusion model implementations, takes a more direct approach that builds upon the principles of Denoising Diffusion Probabilistic Models (DDPMs) [18, 27]. We train a single neural network to directly approximate the full conditional score, $s_{\theta}(x_{t},y,t)\approx\nabla_{x}\log p_{t}(x_{t}|y)$. The conditioning information $y$ (the PINO output) is provided to the network as an additional input, typically by concatenating the condition $y$ with the noisy state $x_{t}$ along the channel dimension. The network is then trained to denoise this composite input.
It is worth emphasizing that the efficacy of the entire framework hinges on a crucial division of labor. The score-based denoising process is inherently more effective at refining high-wavenumber components than low-wavenumber ones, where the score function gradient is weaker. Consequently, for the framework to succeed, the neural operator must provide an accurate low-frequency prior, or ‘mean flow’. A detailed analysis supporting this frequency-dependent behavior of the score function is presented in the appendix of Oommen et al. [24].
@@ 2.3 Governing Equations and Data Generation
subject to the constraints $\nabla\cdot\mathbf{u}=0$ and $\nabla\cdot\mathbf{B}=0$. Here, $\mathbf{u}$ is the fluid velocity, $\mathbf{B}$ is the magnetic field, $p$ is the pressure, $\rho_{0}=1$ is the constant fluid density, $\nu$ is the kinematic viscosity, and $\eta$ is the magnetic diffusivity. The kinetic and magnetic Reynolds numbers are defined as $\mathrm{Re}=1/\nu$ and $\mathrm{Re_{m}}=1/\eta$, respectively. For this work, we consider a magnetic Prandtl number $\mathrm{Pr_{m}}=\nu/\eta=1$, such that $\mathrm{Re}=\mathrm{Re_{m}}$. We deliberately restrict our attention to the two-dimensional, incompressible, resistive MHD equations at a unit magnetic Prandtl number. This controlled regime provides a stringent yet interpretable testbed, retaining essential turbulent features such as nonlinear cascades and magnetic reconnection, while allowing for reproducible dataset generation and systematic validation of physics-informed AI surrogates. We note, however, that the realm of applicability of the present results does not extend to fully three-dimensional turbulence, where anisotropy, Alfvénic dynamics, and small-scale intermittency introduce qualitatively new challenges. Likewise, regimes with magnetic Prandtl numbers far from unity—which are relevant for both astrophysical and laboratory plasmas—exhibit distinct dissipative scale separations and dynamo behaviors that are not addressed here. These extensions require substantially more work for dataset production and careful hyperparameter tuning for stable training, and are the subject of ongoing work. The aim of this manuscript is therefore focused: to establish, within this well-defined 2D Prandtl number $\mathrm{Pr_{m}}=1$ setting, that turbulent magnetized flows can be effectively modeled using physics-informed neural operators augmented by diffusion-based generative corrections.
@@
This system of equations was solved numerically using the open-source spectral solver Dedalus [29], closely following the numerical framework established in Ref. [17]. The simulations were configured on a doubly periodic 2D domain $0,1]^{2}$, which was discretized using a Fourier basis with a primary resolution of $128\times 128$ grid points and a de-aliasing factor of $3/2$. A fourth-order Runge-Kutta scheme was employed for time integration with a fixed timestep of $\Delta t=10^{-3}$, evolving the system to a final time of $t=1.0$ with snapshots saved every $0.01$ time units. This process yields 101 high-resolution temporal frames for each simulation, which were then temporally subsampled by a factor of 4 to produce the final 26-frame sequences used for training and evaluation. To create a diverse ensemble of scenarios, initial conditions were generated from Gaussian Random Fields (GRFs) [30] with a characteristic length scale of $l=0.1$, ensuring the initial velocity and magnetic fields were divergence-free. To investigate the transition from laminar to turbulent flow, we generated distinct datasets for a wide range of Reynolds numbers, specifically $\mathrm{Re}\in\{100,250,500,750,1000,3000,10000\}$. The generation of this extensive dataset was performed on the NCSA Delta cluster, where we employed a massively parallel approach to distribute thousands of independent simulation instances using all 128 CPU cores of a compute node. For each Reynolds number, this process yielded an ensemble of 1000 simulations, which was randomly partitioned into a training set of 800 samples, a validation set of 100 samples, and a test set of 100 samples.
@@ 2.4 Diffusion-Enhanced Operator Framework
The first stage of our pipeline employs a Physics-Informed Tensor-Factorized Fourier Neural Operator (PINO) [14, 11] to learn the mean-field evolution of the MHD system. The architecture, an extension of the model used in Ref. [17], is designed to process 3D spatio-temporal data ($x,y,t$). It consists of 8 Fourier layers with 32 latent channels. Within each spectral convolution, we retain 8 Fourier modes and utilize a Canonical Polyadic (CP) tensor decomposition with a rank of 0.5 to factorize the spectral weights [31, 32]. This factorization enhances model expressivity while maintaining computational tractability on large-scale hardware. Crucially, the PINO is trained with a composite loss function that, in addition to a standard data-fitting term, includes residuals of the governing MHD equations (Eqs. 9-11) as soft constraints. The PINO was trained on a single AMD MI250x GPU on the Frontier supercomputer at the Oak Ridge Leadership Computing Facility (OLCF).
The output of the trained PINO, which represents a smooth, spectrally-limited approximation of the true solution, serves as a conditional prior for the second stage: a score-based diffusion model [19]. This generative model is built upon a UNet backbone [33] with a base dimension of 128 and six down/up-sampling stages, defined by channel multipliers of (1, 2, 3, 5, 8, 12). Self-attention, accelerated via Flash Attention [34, 35], is incorporated at each resolution with 8 heads and a head dimension of 64. The network operates within a preconditioned diffusion framework that uses calibrated noise schedules and input normalizations to balance the learning dynamics. This design has been shown to enhance training efficiency and sample fidelity in recent diffusion models [36]. The model is tasked with stochastically generating the high-wavenumber turbulent structures absent from the PINO’s prediction. During its reverse-denoising process, it is explicitly conditioned on the PINO output, ensuring that the generated turbulent structures are physically consistent with the underlying large-scale flow. The training data for the diffusion model, consisting of pairs of PINO predictions and their corresponding ground truth simulation states, was generated on Frontier, while the model itself was trained on NVIDIA H100 GPUs on the NCSA Delta AI cluster.
@@ 3.2 Spectral Correction via Diffusion-Enhanced Framework
The generative capability of this corrector stage arises from training via denoising score matching [37], where the model learns to approximate the score function, $\nabla_{x}\log p(x;\sigma)$, across a range of noise scales $\sigma$. Inference proceeds by numerically solving a reverse-time stochastic differential equation (SDE), gradually transforming an initial noise sample into a realistic output by following the learned score field. The addition of controlled stochasticity at each step mitigates error accumulation and enhances sample quality [36].
@@ 3.4 Performance at High Reynolds Numbers
3.5 Generalization Across Reynolds Numbers
Our methodology relies on training specialized models for each distinct physical regime. To empirically validate this choice and explore the framework’s generalization capabilities, we conducted a series of cross-regime experiments. First, we assessed performance across a large parametric gap by applying our trained $\mathrm{Re}=1000$ model to the $\mathrm{Re}=100$ test set. The results confirmed a fundamental failure in generalization, with the time-averaged relative L2 error exceeding 0.85. This is attributed to the critical mismatch in learned physical priors: the model trained on turbulent statistics attempted to impose unphysical, high-frequency dynamics onto the smooth, laminar flow. This finding provides strong empirical justification for the necessity of regime-specific models when bridging disparate physical systems.
Next, we investigated generalization to a physically similar regime by applying the same $\mathrm{Re}=1000$ model to a dedicated test set of 100 simulations generated at $\mathrm{Re}=900$, following the same methodology described in Section 2.3. While the PINO-only model still yielded a substantial L2 error of approximately 0.40, demonstrating the sensitivity of the deterministic operator to physical parameters, the full PINO+Diffusion framework proved remarkably robust. The diffusion model, conditioned on the imperfect PINO prediction, successfully corrected for the parametric mismatch, reducing the final L2 error to a promising $\sim$0.13. This result highlights the powerful corrective and adaptive capability of the generative diffusion stage.
These findings collectively suggest that while our framework cannot bridge disparate physical regimes, it exhibits robust potential for interpolation and adaptation between nearby Reynolds numbers. This opens a promising avenue for future work, where a pre-trained model could serve as a powerful foundation, being efficiently adapted to new physical parameters via fine-tuning of the diffusion component, thereby reducing the computational cost of developing new surrogate models.
@@ 4 Summary and Conclusions
In this work, we addressed the challenge of constructing accurate surrogate models for two-dimensional, incompressible MHD, particularly in regimes characterized by complex, multi-scale turbulence. While PINOs are effective in capturing large-scale, equation-constrained dynamics, we demonstrated that they are fundamentally limited by an inherent spectral bias that suppresses high-wavenumber content [17]. This limitation becomes especially pronounced at moderate to high Reynolds numbers, where fine-scale magnetic and velocity structures dominate the system evolution.
@@
These findings establish a promising paradigm for surrogate modeling in computational physics. By delegating the learning of low-frequency, deterministic evolution to a neural operator and the high-frequency, stochastic corrections to a generative model, the proposed framework provides a principled way to overcome the spectral limitations of existing approaches. Our cross-regime analysis further validates this paradigm, demonstrating not only the necessity of specialized models for disparate physical regimes but also the framework’s robustness to small parametric perturbations. This divide-and-conquer strategy is likely to generalize to a broad class of multi-scale problems in fluid dynamics, astrophysics, and plasma physics.
We conclude by emphasizing that our analysis has been confined to the two-dimensional, incompressible, resistive MHD system at $\mathrm{Pr_{m}}=1$. This deliberate choice enabled controlled dataset generation and systematic evaluation of the DINOs framework while retaining essential turbulent phenomena such as cascades and reconnection. The extension of this approach to fully three-dimensional turbulence—where anisotropy, Alfvénic interactions, and small-scale intermittency become central—as well as to regimes with magnetic Prandtl numbers significantly different from unity, are active directions of our ongoing research. These forthcoming studies will assess the broader applicability of DINOs across the diverse parameter regimes relevant to plasma physics and astrophysics.
Building upon our generalization study, a key direction for future work is the exploration of fine-tuning pre-trained models as a highly efficient strategy for adapting the framework to new, nearby physical parameters. Future work may also explore infusing physics-based constraints—such as PDE residuals—into the diffusion stage to enhance physical consistency [20], as well as extending the framework to kinetic plasma models. These directions offer compelling opportunities to further improve fidelity, generalizability, and interpretability in the data-driven modeling of complex physical systems.
@@ Acknowledgements
EAH acknowledges the partial support of NSF grants OAC-2514142 and OAC-2209892. RH acknowledges partial support from NSF grants OAC-2004879, OAC-2005572, OAC-2103680, OAC-2310548, OAC-2411068. An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used supporting resources at the Argonne and the Oak Ridge Leadership Computing Facilities. The Argonne Leadership Computing Facility at Argonne National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC02-06CH11357. The Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory is supported by the Office of Science of the U.S. DOE under Contract No. DE-AC05-00OR22725. This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
@@
The complete source code for the Diffusion-Integrated Neural Operators (DINOs) framework developed for this study, including all modules for data generation, model training, and analysis, is publicly available in the official repository at https://github.com/semihkacmaz/DINOs. Additionally, the trained models are deposited in the Garden, a platform for sharing and running scientific AI models, and are available at the following address: https://thegardens.ai/#/garden/10.26311%2Fc4bj-8h61.
Appendix A: Detailed Results for the High-Turbulence Regime